OpenStudy (anonymous):

solve the system by the elimination method 5x+5y=-7 7x-3y=17

OpenStudy (anonymous):

5x+5y=-7 [multiply by 3] gives 15x+15y=-21 [a] 7x-3y=17 [multiply by 5] gives 35x+15y=-85 [b] Subtract [a] from [b] 20x = -64 [divide by 20] x = -3.2 substitute this into 5x + 5y = -7 5[-3.2] + 5y = -7 -16 + 5y = -7 [add 16 to both sides] 5y = 9 y = 9/5